# Work statistics for Quantum Spin Chains: characterizing quantum phase transitions, benchmarking time evolution, and examining passivity of quantum states

### v3

@@ VI Numerical results for examining the passivity of thermal/ground states
Figure 8: Demonstration of passivity of thermal states in $n=10$ quantum Ising chain under impulse process implemented by the Hamiltonian of (11) and (13) with $H_{0}$ of (31) with $h_{x}=h_{z}=0$ and with $V=\sum_{k}S_{k}^{x}$, $\sum_{k}S_{k}^{y}$ or $\sum_{k}S_{k}^{z}$. The resultant average work done per $\overline{w}^{\rm impulse}$ for $\beta=1$ are obtained by ED and TEBD. Its positiveness for all considered $V$’s and $\lambda\in[0,0.5]$ demonstrates the passivity of thermal states as pointed out in (20).
We now can consider the passivity of a given initial state undergoing a (cyclic) impulse process. Two natural candidate categories are thermal or ground states. The passivity of (relativistic) thermal states was shown to be ensured [35, 36, 37]. For the thermal states of the non-relativistic systems, such as the quantum spin chains, as argued in (20), the passivity should be guaranteed by the fluctuation theorem, i.e., the second law of thermodynamics. Here, we demonstrate this is the case for the $n=10$ Ising chain. The result is shown in Fig. 8, from which we see that the average work done $\overline{w}^{\rm impluse}=-w^{\rm impulse}_{\rm ext}$ for the thermal state at $\beta=1$ is always positive for various $V$ and $\lambda$.
We now turn to the issue of passivity for the ground states under the cyclic impulse processes. As shown in section II.4, the passivity of ground states is guaranteed to be passive under the hermitian action, i.e., $V^{\dagger}=V$ by the variational principle. This is because the average work extraction from a cyclic process can be understood as the energy difference between the ground state and the excited state driven by $V$ so that it is always negative as the variational principle guarantees. On the other hand, there is no such simple interpretation for the average work extraction if the action $V$ is non-hermitian, i.e., one cannot apply the variational principle to check the passivity of ground states. For the latter cases, it is then interesting to examine the passivity by evaluating average work extraction numerically.
As a startup, we numerically examine the passivity of ground states for various hermitian or non-hermitian actions but small $\lambda$, i.e., $\lambda=0.1$. In Fig. 9, We plot the average work done per site $\overline{w}^{\rm impluse}=-w^{\rm impulse}_{\rm ext}$ as the function of $h_{z}$, which labels the ground states of the Ising-like chain under various hermitian and non-hermitian actions. A similar result for the Haldane-like chains is also given in Fig. 13 of Appendix B. To obtain these results, we first evaluate $G(-is)$ by the MPS-TEBD method, and then from it extract $\overline{w}^{\rm impluse}=-w^{\rm impulse}_{\rm ext}$ from $G(-is)$. The ground states considered are all passive under the various hermitian and non-hermitian actions described in Fig. 9. This is a consistency check for our numerical method with the discussion in section II.4 for examining the passivity of ground states of Ising-like chains.
@@
Figure 9: Plot of $\overline{w}^{\rm impulse}$ to show the passivity of ground states of $n=100$ quantum Ising chain with $h_{x}=0$ but varying $h_{z}\in[0,1]$ under impulse process implemented by Hamiltonian of (11) and (13) with $\lambda=0.1$ and hermitian actions $V=\sum_{k}S_{k}^{x}$ (red circles ), $\sum_{k}S_{k}^{y}$ (black cross), $\sum_{k}S_{k}^{z}$ (green cross), and non-hermitian actions $V=\sum_{k}iS_{k}^{x}$ (blue diamonds), $\sum_{k}iS_{k}^{y}$ (yellow square), $\sum_{k}iS_{k}^{z}$ (pink triangles). Again, we see that $\overline{w}^{\rm impulse}$ can also indicate quantum phase transition by sudden jump when compared with the phase diagram of the entanglement entropy.
@@
Figure 10: Plot of $\overline{w}^{\rm impulse}$ to show the passivity of ground states and its pattern as a function of the coupling $\lambda$ of $n=100$ quantum Ising chain. We fix $h_{x}=0$ but vary $h_{z}\in[0,1]$ to obtain various ground states, and also vary $\lambda\in[-10,10]$ under impulse process implemented by Hamiltonian of (11) and (13) with the hermitian actions (a) $V=\sum_{k}S_{k}^{x}$, (b) $V=\sum_{k}S_{k}^{y}$, (c) $V=\sum_{k}S_{k}^{z}$ in the top row, and the non-hermitian actions (d) $V=i\sum_{k}S_{k}^{x}$, (e) $V=i\sum_{k}S_{k}^{y}$ and (f) $V=i\sum_{k}S_{k}^{z}$ in the bottom row.
@@
By exploiting the numerical method’s power in checking the ground states’ passivity, we now consider the cases with quite an extensive range of $\lambda\in[-10,10]$ for both hermitian and non-hermitian actions. The results for the Ising-like chains are shown in Fig. 10, and the similar results for the Haldane-like chains and for the more general non-hermitian actions in the Ising-like chains are shown in Fig. 14 and Fig. 15 of Appendix B, respectively. Since the ones shown in the Appendix B bear similar features as in Fig. 10, we will not discuss them in the main text. The top row of Fig. 10 shows the average work done per site on the ground states labeled by the values of $h_{z}\in[0,1]$ for the hermitian actions $V=\sum_{k}S_{k}^{x,y,z}$, and the bottom row shows the results for the non-hermitian actions $V=i\sum_{k}S_{k}^{x,y,z}$. For the hermitian cases, we see a quite interesting feature that $\overline{w}^{\rm impulse}$ is periodic with respect to the coupling $\lambda$ with the period equal to $2\pi$ or $\pi$. This periodic behavior has been discussed in section II.4 due to the peculiar feature of the impulse process. We will not see such behavior for the cyclic process with a finite time duration. For the non-hermitian case, we first see that all the ground states considered are passive as in the hermitian cases. However, $\overline{w}^{\rm impulse}$ is no longer periodic with respect to $\lambda$ nut levels off as $\lambda$ becomes large. In both cases, the average work done is bounded no matter how large the coupling $\lambda$ is. Again, this is due to the zero time duration of the impulse process, so the work done is limited to such an infinitesimal time duration. However, it is unclear why the passivity remains intact for the non-hermitian action, especially without the guarantee from the variational principle. This issue deserves future study to explore.
@@ VII Conclusions
In this paper, we apply the numerical method for real-time evolution, such as Exact Diagonalization and MPS-based TEBD, to quantum spin lattice models to study the work statistics. We focus on three aspects: (i) show that the average work done can be used as the order parameter for the (quantum) phase transitions, including the topological phase transitions; (ii) we propose to adopt Jarzynski equality as the benchmark for the accuracy of the numerical real-time evolution methods; (iii) we examine the passivity of thermal states and ground states of quantum spin chains under some cyclic impulse processes. Our numerical results show that all the ground states are passive under both hermitian and non-hermitian actions considered in this work. Although the variational principle ensures the passivity of ground states under hermitian actions, it is not the case for non-hermitian actions. It is interesting to explore the reason for the passivity of ground states under non-hermitian action seen in this paper, and also explore the possibility of active ground states by more general actions. Once the active ground states exist, we may adopt them to implement the quantum engine naturally to extract quantum work in the cyclic processes.
@@ Acknowledgements.
We thank Masahiro Hotta and Jhh-Jing (Arthur) Hong for their helpful discussions. FLL is supported by the Taiwan Ministry of Science and Technology through Grant No. 109-2112-M- 003-007-MY3. CYH is supported by the Taiwan Ministry of Science and Technology through Grant No. 112-2112-M-029 -006.
Appendix A Two numerical consistency checks
@@
A.1 Equivalence between thermal representations by MPO and METTS
@@
A.2 Equivalence between average works done obtained from Eq. (10) or from Eq. (16)
Figure 12: Agreement of the average work done from two methods: (i) extraction from the work characteristic function $G(-is)$ of Eq. (10) (black crosses), and (ii) vacuum expectation value of energy offset by Eq. (16) (blue diamonds) for the quantum Ising chain for various $h_{z}$ in the quantum quench process.
@@
There are two ways to obtain the average work done by a sudden quench process. One is to extract it from the work characteristic function $G(u)$ of (10) by using (7). The second way is to calculate it by evaluating the vacuum expectation value of the Hamiltonian offset using (16). In Fig. 12 for $n=10$ quantum Ising model for various $h_{z}$ in the initial Hamiltonian $H_{0}$, we show that the numerical results from both methods agree. This serves as a consistency check for the TEBD method.
Appendix B More results for examining the passivity of ground states in Ising-like and Haldane-like chains
Figure 13: Plot of $\overline{w}^{\rm impulse}$ to show the passivity of ground states of $n=100$ Haldane-like chain with $J_{z}=1.5$ but varying $D\in[0,2]$ under impluse process of $\lambda=0.1$ and hermitian actions $V=\sum_{k}S_{k}^{x}$ (black triangles), $\sum_{k}S_{k}^{y}$ (yellow right-arrows), $\sum_{k}S_{k}^{z}$ (red circles), $\sum_{k}(S_{k}^{x})^{2}$ (green diamonds), $\sum_{k}(S_{k}^{y})^{2}$ (orange left-arrows) and $\sum_{k}(S_{k}^{z})^{2}$ (blue squares). The $\overline{w}^{\rm impluse}$ plots show that the ground states are always passive for all considered $V$’s. The degeneracy of $\overline{w}^{\rm impluse}$ under swapping of $S_{i}^{x}$ and $S_{i}^{y}$, as shown here, can be understood as the $Z_{2}$ symmetry of $H_{0}$ and $SO(3)$ spin algebra under the transformation: $(S_{k}^{x},S_{k}^{y},S_{k}^{z})\rightarrow(-S_{k}^{y},S_{k}^{x},S_{k}^{z})$. Besides, $\overline{w}^{\rm impluse}$ always vanishes for $V=\sum_{k}S_{k}^{z}$ because of $[H_{0},\sum_{k}S_{k}^{z}]=0$.
Figure 14: Plot of $\overline{w}^{\rm impulse}$ to show the passivity of ground states and its pattern as a function of the coupling $\lambda$ of $n=100$ Haldane-like chain with $J_{z}=1.5$. We vary the coupling $D\in[0,2]$ to obtain various ground states, and also vary $\lambda\in[-10,10]$ under impulse process implemented by Hamiltonian of (11) and (13) for the hermitian actions (a) $V=\sum_{k}S_{k}^{x}$, (b) $V=\sum_{k}(S_{k}^{x})^{2}$, (c) $V=\sum_{k}(S_{k}^{z})^{2}$ in the top row, and non-hermitian actions (d) $V=i\sum_{k}S_{k}^{x}$, (e) $V=\sum_{k}(S_{k}^{x}+iS_{k}^{y})$, and (f) $V=i\sum_{k}(S_{k}^{x})^{2}$ in the bottom row.
@@ Appendix B More results for examining the passivity of ground states in Ising-like and Haldane-like chains
Figure 15: Plot of $\overline{w}^{\rm impulse}$ to show the passivity of ground states and its pattern as a function of the coupling $\lambda$ of $n=100$ quantum Ising chain. We fix $h_{x}=0$ but vary $h_{z}\in[0,1]$ to obtain various ground states, and also vary $\lambda\in[-10,10]$ under impulse process implemented by Hamiltonian of (11) and (13) with the non-hermitian actions (a) $V=\sum_{k}(S_{k}^{x}+iS_{k}^{y})$, (b) $V=\sum_{k}(S_{k}^{y}+S_{k}^{y})$,and (c) $V=\sum_{k}(S_{k}^{x}+S_{k}^{z})$.

### v4

@@ VI Numerical results for examining the passivity of thermal/ground states
Figure 8: Demonstration of passivity of thermal states in $n=10$ quantum Ising chain under impulse process implemented by the Hamiltonian of Eq. (11) and Eq. (13) with $H_{0}$ of Eq. (33) with $h_{x}=h_{z}=0$ and with $V=\sum_{k}S_{k}^{x}$, $\sum_{k}S_{k}^{y}$ or $\sum_{k}S_{k}^{z}$. The resultant average work done per site $\overline{w}^{\rm impulse}$ for $\beta=1$ are obtained by ED and TEBD. Its positiveness for all considered $V$’s and $\lambda\in[0,0.5]$ demonstrates the passivity of thermal states as pointed out in Eq. (22).
@@
Figure 9: Plot of $\overline{w}^{\rm impulse}$ to show the passivity of ground states of $n=1024$ quantum Ising chain with $h_{x}=0$ but varying $h_{z}\in[0,1]$ under impulse process implemented by Hamiltonian of Eq. (11) and Eq. (13) with $\lambda=0.1$ and Hermitian actions $V=\sum_{k}S_{k}^{x}$ (red circles ), $\sum_{k}S_{k}^{y}$ (black cross), $\sum_{k}S_{k}^{z}$ (green cross), and non-Hermitian actions $V=\sum_{k}iS_{k}^{x}$ (blue diamonds), $\sum_{k}iS_{k}^{y}$ (yellow square), $\sum_{k}iS_{k}^{z}$ (pink triangles). Again, we see that $\overline{W}^{\rm impulse}$ can also indicate quantum phase transition by mild crossover behaviors when compared with the phase diagram of the entanglement entropy of initial state (gray circle).
@@
Figure 10: Plot of $\overline{w}^{\rm impulse}$ to show the passivity of ground states and its pattern as a function of the coupling $\lambda$ of $n=1024$ quantum Ising chain. We fix $h_{x}=0$ but vary $h_{z}\in[0,1]$ to obtain various ground states, and also vary $\lambda\in[-10,10]$ under impulse process implemented by Hamiltonian of Eq. (11) and Eq. (13) with the Hermitian actions (a) $V=\sum_{k}S_{k}^{x}$, (b) $V=\sum_{k}S_{k}^{y}$, (c) $V=\sum_{k}S_{k}^{z}$ in the top row, and the non-Hermitian actions (d) $V=i\sum_{k}S_{k}^{x}$, (e) $V=i\sum_{k}S_{k}^{y}$ and (f) $V=i\sum_{k}S_{k}^{z}$ in the bottom row.
We now can consider the passivity of a given initial state undergoing a (cyclic) impulse process. Two natural candidate categories are thermal or ground states. The passivity of (relativistic) thermal states was shown to be ensured [45, 46, 47]. For the thermal states of the non-relativistic systems, such as the quantum spin chains, as argued in Eq. (22), the passivity should be guaranteed by the fluctuation theorem, i.e., the second law of thermodynamics. Here, we demonstrate this is the case for the $n=10$ Ising chain. The result is shown in Fig. 8, from which we see that the average work done per site $\overline{w}^{\rm impluse}=-w^{\rm impulse}_{\rm ext}$ for the thermal state at $\beta=1$ is always positive for various $V$ and $\lambda$.
We now turn to the issue of passivity for the ground states under the cyclic impulse processes. As shown in section II.4, the passivity of ground states is guaranteed to be passive under the Hermitian action, i.e., $V^{\dagger}=V$ by the variational principle. This is because the average work extraction from a cyclic process can be understood as the energy difference between the ground state and the excited state driven by $V$. It is always negative as the variational principle guarantees. On the other hand, there is no such simple interpretation for the average work extraction if the action $V$ is non-Hermitian, i.e., one cannot apply the variational principle to check the passivity of ground states. For the latter cases, it is then interesting to examine the passivity by evaluating average work extraction numerically.
As a startup, we numerically examine the passivity of ground states for various Hermitian or non-Hermitian actions but small $\lambda$, i.e., $\lambda=0.1$. In Fig. 9, We plot the average work done per site $\overline{w}^{\rm impluse}=-w^{\rm impulse}_{\rm ext}$ as the function of $h_{z}$, which labels the ground states of the Ising-like chain under various Hermitian and non-Hermitian actions. A similar result for the Haldane-like chains is also given in Fig. 16 of Appendix C. To obtain these results, we first evaluate $G(-is)$ by the MPS-TEBD method, and then from it extract $\overline{w}^{\rm impluse}=-w^{\rm impulse}_{\rm ext}$ from $G(-is)$. The ground states considered are all passive under the various Hermitian and non-Hermitian actions described in Fig. 9. This is a consistency check for our numerical method with the discussion in section II.4 for examining the passivity of ground states of Ising-like chains.
@@
By exploiting the numerical method’s power in checking the ground states’ passivity, we now consider the cases with quite an extensive range of $\lambda\in[-10,10]$ for both Hermitian and non-Hermitian actions. The results for the Ising-like chains are shown in Fig. 10, and the similar results for the Haldane-like chains and the more general non-Hermitian actions in the Ising-like chains are shown in Fig. 17 and Fig. 15 of Appendix C, respectively. Since the ones shown in the Appendix C bear similar features as in Fig. 10, we will not discuss them in the main text. The top row of Fig. 10 shows the average work done per site on the ground states labeled by the values of $h_{z}\in[0,1]$ for the Hermitian actions $V=\sum_{k}S_{k}^{x,y,z}$, and the bottom row shows the results for the non-Hermitian actions $V=i\sum_{k}S_{k}^{x,y,z}$. For the Hermitian cases, we see a quite interesting feature that $\overline{W}^{\rm impulse}$ is periodic with respect to the coupling $\lambda$ with the period equal to $2\pi$ or $\pi$. This periodic behavior has been discussed in section II.4 due to the peculiar feature of the impulse process. We will not see such behavior for the cyclic process with a finite time duration. For the non-Hermitian case, we first see that all the ground states considered are passive as in the Hermitian cases. However, $\overline{W}^{\rm impulse}$ is no longer periodic with respect to $\lambda$ nut levels off as $\lambda$ becomes large. In both cases, the average work done is bounded no matter how large the coupling $\lambda$ is. Again, this is due to the zero time duration of the impulse process, so the average work done is limited to such a tiny time duration. However, it is unclear why the passivity remains intact for non-Hermitian action, especially without the guarantee by the variational principle. This issue deserves future study to explore.
@@ VII Conclusions
In this paper, we apply the numerical method for real-time evolution, such as Exact Diagonalization (ED) and MPS-based TEBD, to quantum spin lattice models to study the work statistics. With the power of MPS formalism, we can evaluate the chains up to $1024$ sites, which can effectively suppress the finite size effect. We focus on three aspects: (i) study the behaviors of the moments and cumulants of the work statistics near the quantum phase transitions and examine their capability to indicate quantum critical points; Our results up to the fourth cumulant show that the work statistics can detect the quantum phase transitions charactered with local order parameters, but just barely for the topological phase transitions; (ii) we propose to adopt Jarzynski equality as the benchmark for the accuracy of the numerical real-time evolution methods; (iii) we examine the passivity of thermal states and ground states of quantum spin chains under some cyclic impulse processes. Our numerical results show that all the ground states are passive under both Hermitian and non-Hermitian actions considered in this work. Although the variational principle ensures the passivity of ground states under Hermitian actions, it is not the case for non-Hermitian actions. It is interesting to explore the reason for the passivity of ground states under non-Hermitian action seen in this paper, and also explore the possibility of active ground states by more general actions. Once the active ground states exist, we may adopt them to implement the quantum engine naturally to extract quantum work in the cyclic processes.
@@ Acknowledgements.
We thank Masahiro Hotta and Jhh-Jing (Arthur) Hong for their helpful discussions. FLL is supported by the Taiwan Ministry of Science and Technology through Grant No. 112-2112-M-003 -006 -MY3. CYH is supported by the Taiwan Ministry of Science and Technology through Grant No. 112-2112-M-029 -006.
Appendix A Tentative study of the density of state in Eq. (21)
In section II.3, we argue why the average work done $\overline{W}$ can be used as the local order parameter for quantum phase transition. As $\overline{W}$ can be expressed in (21) as the average over the eigen-energies of the final Hamiltonian, it will have a sudden change when crossing the quantum critical point because the probabilities of states $p(f)$ in (21) for the gapped and the gapless phases have pretty different behaviors at low energy of an extensive system.
Therefore, it is interesting to calculate $p(f)$ and examine the behaviors of the gapped and gapless phases of an extensive spin chain. Unfortunately, the MPS-based method used for the main calculations of the long spin chain can not applied to the excited states. We can only study $p(f)$ tentatively by using ED method for the short spin chains. In this case, we will miss the ground state degeneracy for the gapless phases due to the sizeable finite-size effect. Despite that, we still present the ED results of $p(f)$ as a function of eigen-energy $E_{f}$ of the $n=10$ quantum spin chains with $h_{z}=0.4$, $0.5$, $1.0$ under a sudden quench process with $V=\sum_{k}S_{k}^{z}$ and $\Delta\lambda=0.1$. The result is shown in Fig. 11. As expected, due to the large finite size effect for $n=10$, it is hard to see the effect of ground state degeneracy on $p(f)$ for either $h_{z}=0.5$ (black crosses) with the initial Hamiltonian in the critical phase or $h_{z}=0.4$ (red circles) with the final Hamiltonian in the critical phase as $\Delta\lambda=0.1$. However, we can see that the $p(f)$’s for these two cases are more focused than the one with $h_{z}=1.0$ (green pluses). This implies the tendency for the effect of ground state degeneracy.
Figure 11: The ED results of weight $p(f)$ from Eq. (21) are plotted as a function of the eigen-energy $E_{f}$ of the Hamitain $H_{+}=H_{0}+\Delta\lambda V$ for the n = 10 quantum Ising chains with $h_{z}=0.4$, $0.5$, $1.0$ under a sudden quench process with $V=\sum_{k}S_{k}^{z}$ and $\Delta\lambda=0.1$. We can see that the $p(f)$’s for $h_{z}=0.5$ (black crosses) and $h_{z}=0.4$ (red circles) have a more focused energy spectrum than the one for $h_{z}=1.0$ (green pluses). Note that for $h_{z}=0.5$ ($h_{z}=0.4$), the initial (final) Hamiltonian is in the critical phase because $\Delta\lambda=0.1$. Thus, the more focused energy spectrum for these two cases implies a tendency for the effect of ground-state degeneracy.
Appendix B Two numerical consistency checks
@@
B.1 Equivalence between thermal representations by MPO and METTS
@@
B.2 Equivalence between average works done obtained from Eq. (10) or from Eq. (16)
@@
There are two ways to obtain the average work done by a sudden quench process. The first one is to extract it from the work characteristic function $G(u)$ of Eq. (10) by using Eq. (7). The second one calculates it by evaluating the vacuum expectation value (vev) of the Hamiltonian offset using Eq. (16). We check the agreement of the results from both ways for the quantum Ising and Haldane chains of $n=10$ by ED or $n=1024$ by MPS-based method, as shown in Fig. 13 and Fig. 14, respectively.
In Fig. 13 for the quantum Ising model with various $h_{z}$ in the initial Hamiltonian $H_{0}$ under sudden quench process with $V=\sum_{k}S_{k}^{z}$ and $\Delta\lambda=0.01$, we show that the numerical agreement of the numerical results of $\overline{W^{m}}$ for $m=1,2,3,4$ obtained from the characteristic function $G(-is)$ (red crosses) or from the vev of $V^{m}$ (blue circles). This serves as a consistency check for the TEBD method.
In Fig. 14 for $n=1024$ Haldane (XXZ) chain with various coupling $D$ in the initial Hamiltonian $H_{0}$ under sudden quench process with $V=\sum_{k}(S_{k}^{z})^{2}$ and $\Delta\lambda=0.01$, we show that results for (a) $\overline{W}$, (b) $\overline{W^{2}}$, (c) $\sigma^{2}_{W}$ and (d) entanglement entropy $S$, which are obtained from the characteristic function $G(-is)$ (black dashed line for $\chi_{cut}=32$ or solid line for $\chi_{cut}=64$) or from the vev of $V^{m}$ (red crosses for $\chi_{cut}=64$ or blue circles $\chi_{cut}=64$). We can see that the results from both ways match well for (a) and (c) but not for (b). This mismatch can be due to the need of large $\chi_{cut}$ to compensate for the errors of performing numerical derivatives on $G(-is)$. This is seen from the improvement of $\overline{W^{2}}$ in (b) and $S$ in (d) when enlarging $\chi_{cut}$ twice large.
Figure 13: Agreement of $\overline{W^{m}}$ of $m=1,2,3,4$ obtained from either characteristic function $G(-is)$ of Eq. (10) (red crosses) or from the vev of $V^{m}$ (blue circles) for the transverse quantum Ising chains under sudden quench process with $V=\sum_{k}S_{k}^{z}$ and $\Delta\lambda=0.01$. In (a) and (b), the result for $n=10$ chains is obtained by ED method, and for (a) it also shows the result from the vev of Hamiltonian offset by Eq. (16) (green squares). In (c),(d),(e), and (f), the results for $n=1024$ chains are calculated by the MPS-based method.
Figure 14: Examining the agreement of (a) $\overline{W}$, (b) $\overline{W^{2}}$, (c) $\sigma^{2}_{W}$ for $n=1024$ Haldane (XXZ) chain with various coupling $D$ in the initial Hamiltonian $H_{0}$ of size $n=1024$ under sudden quench process with $V=\sum_{k}(S_{k}^{z})^{2}$ and $\Delta\lambda=0.01$. These results are obtained from the characteristic function $G(-is)$ (black dashed line for $\chi_{cut}=32$ or solid line for $\chi_{cut}=64$) or from the vev of $V^{m}$ (red crosses for $\chi_{cut}=64$ or blue circles $\chi_{cut}=64$). We see that the results from both ways match well for (a) and (c) but not for (b). We thus evaluate entanglement entropy $S$ in (d) to show the relevance of increasing the bond dimension $\chi_{cut}$ in improving the accuracy, which could also be the reason for the mismatch in (b).
@@
Figure 15: Plot of $\overline{w}^{\rm impulse}$ to show the passivity of ground states and its pattern as a function of the coupling $\lambda$ of $n=1024$ quantum Ising chain. We fix $h_{x}=0$ but vary $h_{z}\in[0,1]$ to obtain various ground states, and also vary $\lambda\in[-10,10]$ under impulse process implemented by Hamiltonian of Eq. (11) and Eq. (13) with the non-Hermitian actions (a) $V=\sum_{k}(S_{k}^{x}+iS_{k}^{y})$, (b) $V=\sum_{k}(S_{k}^{y}+S_{k}^{y})$,and (c) $V=\sum_{k}(S_{k}^{x}+S_{k}^{z})$.
Appendix C More results for examining the passivity of ground states in Ising-like and Haldane-like chains
In this Appendix, we present more results of examining the passivity of the ground states. The first one is shown in Fig. 15 for the plot of $\overline{W}^{\rm impulse}$ of Ising-like chains as a function of $\lambda\in[-10,10]$ for the more general non-Hermitian actions. It can be seen as the non-Hermitian counterpart of Fig. 10.